Casey Motors recently reported the following information: ∙

Net income = $875,000. ∙

Tax rate = 40%. ∙

Interest expense = $200,000. ∙

Total invested capital employed = $9 million. ∙

After-tax cost of capital = 10%.

What is the company’s EVA?
Answer #1

Earnings before tax = Net income / ( 1- tax rate)

= 875,000/ ( 1-40%)

= 1,458,333.33

Earnings before interest and tax = 1,458,333.33 + 200,000 = 1,658,333.33

EVA = EBIT * (1- Tax ) - Capital employed * Cost of capital

= 1,658,333.33 * (1-40%) - 9,000,000 *10%

= 95,000